OpenStudy (connie21):

Explain why Cos (-330 degrees) is or not equal to Cos (390 degrees).

OpenStudy (anonymous):

They are both the same, 390=360+30, -330=-360+30 and \[\cos(30)=\sqrt{3}/2.\] Your calculator just might not know it.